OpenStudy (kobeni-chan):

Find the sum of a finite geometric sequence from n = 1 to n = 6, using the expression −2(5)^(n − 1).

OpenStudy (kobeni-chan):

@Nnesha

OpenStudy (kobeni-chan):

I know you're supposed to use Sn= a1 - an(r) / 1-r, and I had a1=-2 and an (or a6) = -6250. But when I tried to solve I didnt get any of my multiple choice answers.

Nnesha (nnesha):

formula is \[\huge\rm s_n = \frac{ a_1(1-r^n )}{ 1-r }\] where r is common ratio n is number of terms

OpenStudy (kobeni-chan):

Oh I must've gotten the wrong formula :( So r is 5 and n is 6, right?

Nnesha (nnesha):

yep

OpenStudy (kobeni-chan):

Oh I got it now. Thank you! :)

Nnesha (nnesha):

np :-)